# DEPF: A UAV Multispectral Object Detector with Dual-Domain Enhancement and Priority-Guided Mamba Fusion

### v1

@@ Abstract
Multispectral remote sensing object detection is one of the important application of unmanned aerial vehicle (UAV). However, it faces three challenges. Firstly, the low-light remote sensing images reduce the complementarity during multi-modality fusion. Secondly, the local small target modeling is interfered with redundant information in the fusion stage easily. Thirdly, due to the quadratic computational complexity, it is hard to apply the transformer-based methods on the UAV platform. To address these limitations, motivated by Mamba with linear complexity, a UAV multispectral object detector with dual-domain enhancement and priority-guided mamba fusion (DEPF) is proposed. Firstly, to enhance low-light remote sensing images, Dual-Domain Enhancement Module (DDE) is designed, which contains Cross-Scale Wavelet Mamba (CSWM) and Fourier Details Recovery block (FDR). CSWM applies cross-scale mamba scanning for the low-frequency components to enhance the global brightness of images, while FDR constructs spectrum recovery network to enhance the frequency spectra features for recovering the texture-details. Secondly, to enhance local target modeling and reduce the impact of redundant information during fusion, Priority-Guided Mamba Fusion Module (PGMF) is designed. PGMF introduces the concept of priority scanning, which starts from local targets features according to the priority scores obtained from modality difference. Experiments on DroneVehicle dataset and VEDAI dataset reports that, DEPF performs well on object detection, comparing with state-of-the-art methods. Our code is available in the supplementary material.
Figure 1: (a) The feature difference between low-light RGB images and infrared images. (b) The comparison between Z-order scanning and our Priority-guided scanning.
Introduction
@@
Unmanned aerial vehicles (UAV) play an important role in civil and public fields, including pest controlling, transportation monitoring and relief work etc. (Zhou et al. 2023) In addition, multispectral object detection is one of the most basic application of UAV, which depends on the remote sensing images provided by RGB camera and infrared imaging sensor. The RGB images contain color and texture-details about the real world, but their performance is limited by the low-light condition or overexposure. Instead, the infrared images provide stable capture of thermal data and target’s outlines under any visual conditions, complementing the advantages offered by RGB images.
@@
Therefore, many studies have focused on designing UAV-based multispectral object detection frameworks. According to the fusion mechanism, these methods could be categorized into pixel-wise fusion and feature-wise fusion.
The pixel-wise fusion retains most part of the nature details about the images. TarDAL (Liu et al. 2022) introduces generative adversarial network to reserve the common feature of two modalities in the fused images. SuperYOLO (Zhang et al. 2023a) adopts input-level fusion of RGB and infrared images, enhancing feature representation by the super-resolution branch, while feature-wise fusion operates on multi-stage backbone features. To align fusion with downstream tasks, DetFusion (Sun et al. 2022a) guides training through detection loss derived from fused images.
@@
Different from pixel-wise fusion, the feature-wise fusion enhance the performance of detectors by leveraging the complementarity of multi-modality features. Halfway Fusion (Liu et al. 2016) is the first work that proposes feature-wise fusion for multispectral pedestrian detection. Later, (Sun et al. 2022b) designs UA-CMDet and a novel RGB-Infrared drone dataset named DroneVehicle. However, UA-CMDet adopts simple feature concatenation to achieve multispectral fusion. (Zhang et al. 2022) introduces complementary image and discriminative feature fusion to make full use of the effective information between the two modalities and filter out redundant features. To address inaccurate cross-modal fusion, $\text{C}^{2}\text{Former}$ (Yuan and Wei 2024) employs CNNs for feature extraction, while utilizes a Transformer-based Inter-modality Cross-Attention module to generate aligned and complementary features. Inspired by the linear complexity of Mamba, DMM (Zhou et al. 2025b) designs disparity-guided mamba fusion to adaptively merge features and achieve lightweight object detection.
@@
However, the above methods also face three challenges: 1) Although the infrared images provide supplementary information about the scenarios. The low-light RGB images reduces the complementarity during multi-modality fusion, as the obvious differences in their feature maps. As shown in Figure 1 (a). 2) In the fusion stage, the local small target modeling are interfered with redundant information. Specially, as shown in Figure 1 (b), the Mamba-based fusion methods like DMM adopts Z-order scanning focuses on global modeling and suppresses the local dependencies of 2D images. Therefore, the feature map contains noise. 3) The transformer-based methods like $\text{C}^{2}\text{Former}$ has large computational costs for detection task.
@@
To address these challenges, a UAV Multispectral Object Detector with Dual-domain Enhancement and Priority-guided Mamba Fusion named DEPF is proposed. Firstly, to enhance low-light remote sensing RGB images, DDE Module is designed, which contains CSWM block and FDR block. Specially, CSWM block contains Cross-Scale Mamba Scanning for the low-frequency (LF) components from 2D discrete wavelet transform (2D-DWT) to enhance the global brightness of images. FDR block constructs Spectrum Recovery Network (SRN) to enhance the frequency spectra features to recover the texture-details. Secondly, PGMF Module is designed to enhance the local target modeling for multispectral fusion. PGMF constructs Priority Score Network (PSN) to obtain the priorities of tokens from feature differences. The priority scores guide serialization to prioritize tokens containing local target features.
@@
    DEPF is proposed to enhance the low-light RGB images and improve complementarity in the fusion stage, as well as enhance the local target modeling and achieve feature alignment during fusion.
@@
    PGMF module is designed to focus on local target modeling, reducing the interference of redundant information. To our best knowledge, PGMF is the first approach to improve Mamba mechanism for multi-modality fusion.
@@
    The experiment results on DroneVehicle dataset and VEDAI dataset show that DEPF performs well on object detection and achieves lightweight computation.
Figure 2: The overall architecture of DEPF. (a) is the pipeline of DEPF, (b) is the architecture of DDE module, while (c) is the processing of CSWM block including the Cross-Scale Scanning Mechanism. (d) shows the Cross-Scale Serialization in the CSWM block.
Related Work
Multispectral Object Detection
@@
Pixel-wise fusion aggregates two spectral data into a whole images ahead of schedule, and feeds the images to the detector. (Cao et al. 2023) proposes a local–global adaptive dynamic learning method to fuse multi-modal data. To ensure the fusion network is task-optimized, DetFusion (Sun et al. 2022a) employs the detection loss computed from the fused images as guidance during training. Employing input-level fusion of RGB and IR images. Feature-wise fusion focuses on the features from different stages of backbone network. These stage features are fused and fed to the detection head. TSFADet (Yuan, Wang, and Wei 2022) designs an alignment module to predict modality deviations with subsequent feature map calibration. (Zhao et al. 2024) propose a coarse-to-fine framework to remove redundant information and select features for fusion dynamically. To resolve modality calibration and fusion inaccuracies, $\text{C}^{2}\text{Former}$ (Yuan and Wei 2024) introduced an inter-modal cross-attention module, while its Adaptive Feature Sampling module mitigated computational costs from global attention.
@@
Low-light Images Enhancement
@@
There are many frameworks having been designed for enhancing the low-light images, which lacks enough texture-details due to under exposed characteristic. Brightness enhancement in FourLLIE (Feng et al. 2022) is achieved by dual-stage estimation of amplitude transformation maps within the frequency domain. DiffLL (Jiang et al. 2023) implements sequential denoising refinements in diffusion models to generate realistic details for low-light image enhancement task, highlighting their practical viability. (Zhou et al. 2025a) proposes a framework with guidance of generative perceptual priors obtained from vision-language models, it shows good generalization on real-world data. WaveMamba (Zou et al. 2024) combines wavelet transformation and Mamba mechanism to achieve low-light images restoration. These works enhance the texture-details of instances in low-light images, which gives important help to the detection task.
Vision Mamba
@@
Mamba (Gu and Dao 2023) mechanism has been paid attention to computer vision field because it can model long-distance dependencies and maintain linear computational complexity. Many vision mamba works has developed and performed well in the vision researches. The visual state-space model VMamba (Liu et al. 2024) preserves SSM’s global receptive field capability but operates at linear computational cost, enabling order-of-magnitude speed gains. Vim (Zhu et al. 2024) proposes a new vision backbone with bidirectional scanning mamba, This architecture embeds positional cues into sequences while compacting representations through bidirectional state-space compression. Vision mamba is widely adopted for images classification tasks, like SpectralMamba (Yao et al. 2024), $\text{S}^{2}\text{Mamba}$ (Wang et al. 2025) etc., and images segmentation tasks like $\text{RS}^{3}\text{Mamba}$ (Ma, Zhang, and Pun 2024), VM-Unet (Ruan, Li, and Xiang 2024) etc. These works could achieve good performance with low computation complexity.
@@
However, most of vision mamba works adopts Z-order serialization for scanning the images, which focuses on global context and suppresses the local dependencies. Therefore, for the remote sensing images with many small objects, the background information brings degradation to local targets modeling in the multi-modality fusion stage under the Z-order serialization.
Method
Dual-Domain Mamba Enhancement Module
@@
The low-frequency component in wavelet-domain contains most of information about low-light RGB images, while the high-frequency component only contains small parts of texture-details (Zou et al. 2024). Therefore, the low-frequency component could be adopted to enhance the overall brightness of images. For the texture-details, it is worth noting that frequency domain has good global modeling capacity. The amplitude spectrum contains the energy of details while the phase spectrum contains the structure of the instances in the images. Therefore, these two spectra could be adopted to recover the texture details of images.
@@
Following the above ideas, DDE module is designed for enhancing the low-light RGB remote sensing images, which is shown in Fig.1. DDE module consists of two parts, the first part is CSWM block, which is designed for low-frequency component enhancement. The second part is FDR block, which is designed to enhance the texture details of images.
Cross-Scale Wavelet Mamba Block
@@
The $LL$ is enhanced by CSWM block, which is designed by adopting the strong context-aware properties of Mamba (Albert Gu and Tri Dao, 2023). To ensure that the local objects dependencies are enhanced, $LL$ is conveyed to three convolution layers with kernel sizes $i\in\left[3,5,7\right]$ to obtain multi-scale feature and global feature:
@@
Fourier Details Recovery
The enhanced images $I_{ve}^{n}$ also lacks texture-details after low-frequency enhancement. Therefore, $I_{ve}^{n}$ is conveyed to FDE module for recovering most of texture-details. FDE contains spectrum recovery networks which could enhance the details of amplitude spectrum (AS) and phase spectrum (PS) obtained by 2D-FFT. The enhanced MS and PS are then transformed into time-domain to obtain rich texture-details images $I_{VE}$. This process could be described as:
@@
where $SRN()$ denotes the spectrum recovery network, which consists of two DWCONV layers, one ReLU layer, one LayerNorm layer and one Linear layer. $I_{VE}$ should be conveyed to CSWM block if it is not in the last level of 2D-DWT.
Priority-Guided Mamba Fusion Module
For the feature maps of remote sensing images, there is an obvious difference in the proportion of local target features and background features. However, the Z-order serialization focuses on global modeling, which brings redundant information to weaken the local targets modeling. Therefore, if the scanning prioritizes the feature of local targets, not only does the impact of redundant information be lessened, but also the original global feature is retained. The fused feature map with obvious targets details could be obtained by scanning the priority sequences of two modalities. In the appendix file, we conduct theoretical proof about the advantage of Priority-guided for Mamba modeling.
Following the above ideas, Priority-Guided Mamba Fusion Module (PGMF) is designed to achieve high effective RGB-IR feature fusion, which is shown in Figure 4.
The input feature $\mathcal{F}_{v}$ and $\mathcal{F}_{i}$ are conveyed to small MLPs to reflect into latent space. Moreover, it is worth noting that not only does the feature difference $\varDelta\mathcal{F}$ contain location information of the targets, but also show the regions that be focused mistakenly. Therefore, $\varDelta\mathcal{F}$ is conveyed to Priority Score Network to align feature difference and obtain priority matrices $PMat$. Both of two modalities could obtain their priority features by adding with $PMat$:
@@
The serialization sets up to obtain the sequences by following the descent priority scores, which separates the target from interference information. Therefore, the scanning starts from the targets and ends up with the background information. Due to the difference of two modality, the priority orders of them may not consistent.
@@
Where $Ser()$ denotes serialization operator and $argSort()$ denotes obtaining the tokens after sorting.
Two sequences are fused together and conveyed to SSM module. the input features of two modalities are added with priority feature by residual connection:
@@
The pseudocode and performance analysis of PGMF module could be seen in the appendix file of the supplementary materials.
Loss Function

### v2

# DEPFusion: Dual-Domain Enhancement and Priority-Guided Mamba Fusion for UAV Multispectral Object Detection

@@ Abstract
Multispectral object detection is an important application for unmanned aerial vehicles (UAVs). However, it faces several challenges. First, low-light RGB images weaken the multispectral fusion due to details loss. Second, the interference information is introduced to local target modeling during multispectral fusion. Third, computational cost poses deployment challenge on UAV platforms, such as transformer-based methods with quadratic complexity. To address these issues, a framework named DEPFusion consisting of two designed modules, Dual-Domain Enhancement (DDE) and Priority-Guided Mamba Fusion (PGMF) , is proposed for UAV multispectral object detection. Firstly, considering the adoption of low-frequency component for global brightness enhancement and frequency spectra features for texture-details recovery, DDE module is designed with Cross-Scale Wavelet Mamba (CSWM) block and Fourier Details Recovery (FDR) block. Secondly, considering guiding the scanning of Mamba from high priority score tokens, which contain local target feature, a novel Priority-Guided Serialization is proposed with theoretical proof. Based on it, PGMF module is designed for multispectral feature fusion, which enhance local modeling and reduce interference information. Experiments on DroneVehicle and VEDAI datasets demonstrate that DEPFusion achieves good performance with state-of-the-art methods.
I Introduction
@@
Unamaned aerial vehicles (UAVs) play an important role in civil and public fields, including pest controlling, transportation monitoring and relief work etc. [1] In addition, object detection is one of the most basic application of UAV, which depends on the remote sensing images provided by RGB camera and infrared imaging sensor. The RGB images contain color and texture-details about the real world, but their performance is limited by the low-light condition or overexposure. Instead, the infrared images provide stable capture of thermal data and target’s outlines under any visual conditions, complementing the advantages offered by RGB images.
@@
Therefore, many studies have focused on designing UAV-based object detection frameworks. According to the chosen sensor modality, these methods could be categorized into RGB-based and multispectral object detection.
The RGB-based object detection method relies on the color and texture-details provided by cameras on the UAVs. Sun et.al [2] modify YOLOv8 [3] which exists an issue of losing small object, obtaining higher accuracy of detection. Xiao et.al [4] propose a lightweight framework for small object detection based on multi-scale feature fusion. Lin et al. [5] propose GE-FSOD to address the challenge of limited labeled data and poor generalization in remote sensing object detection. Liang et al. [6] propose the FS-SSD model with spatial context analysis to tackle small object detection in UAV imagery. Yu et al. [7] propose ARPN to address the challenges of detecting tiny persons in UAV imagery, where targets are small in complex backgrounds. Overall, these works advance to RGB-based object detection for UAVs application.
@@
The multispectral object detection introduces infrared images as auxiliary branch to overcome the low-light issue of RGB images. TarDAL [8] introduces generative adversarial network to reserve the common feature of two modalities in the fused images. SuperYOLO [9] adopts input-level fusion of RGB and infrared images, enhancing feature representation by the super-resolution branch, while feature-wise fusion operates on multi-stage backbone features. To align fusion with downstream tasks, DetFusion [10] guides training through detection loss derived from fused images. Sun et.al [11] design UA-CMDet and a novel RGB-Infrared drone dataset named DroneVehicle. However, UA-CMDet adopts simple feature concatenation to achieve multispectral fusion. Zhang et.al [12] introduce complementary image and discriminative feature fusion to make full use of the effective information between the two modalities and filter out redundant features. To address inaccurate cross-modal fusion, $\text{C}^{2}\text{Former}$ [13] employs CNNs for feature extraction, while utilizes a Transformer-based Inter-modality Cross-Attention module to generate aligned and complementary features. Inspired by the linear complexity of Mamba, DMM [14] designs disparity-guided mamba fusion to adaptively merge features and achieve lightweight object detection.
Figure 1: The modality feature difference and the proposed priority-guided scanning strategy. (a) The feature difference between low-light RGB images and infrared images. (b) The comparison between Z-order scanning and our Priority-guided scanning.
@@
However, the above methods also face three challenges: 1) Although the infrared images provide supplementary information about scenarios, the low-light RGB images weaken multispectral fusion due to details loss, causing the obvious differences in their feature maps. As shown in Fig 1 (a). 2) In the fusion stage, the local small target modeling are interfered with redundant information. Specially, as shown in Fig 1 (b), the Mamba-based fusion methods like DMM adopts Z-order scanning focuses on global modeling and suppresses the local dependencies of 2D images. Therefore, the feature map contains noise. 3) The transformer-based methods like $\text{C}^{2}\text{Former}$ has large computational costs for detection task.
@@
To address these challenges, DEPFusion is proposed with Dual-domain Enhancement and Priority-guided Mamba Fusion for multispectral object detection. Firstly, to enhance low-light RGB images, DDE Module is designed, which contains CSWM and FDR block. Specially, CSWM block contains Cross-Scale Mamba Scanning for the low-frequency components from 2D discrete wavelet transform (2D-DWT) to enhance the global brightness of images. FDR block constructs Spectrum Recovery Network (SRN) to enhance the frequency spectra features to recover the texture-details. Secondly, Priority-Guided Serialization is proposed with theoretical proof, which adopting priority scores to guide the scanning of Mamba to prioritize tokens containing local target features. Based on Priority-Guided Serialization, PGMF module is designed to enhance the local target modeling for multispectral fusion. PGMF constructs Priority Score Network (PSN) to obtain the priorities of tokens from feature differences. Experiment results on DroneVehicle and VEDAI dataset demonstrate that DEPFusion achieves lightweight computation and performs well on object detection.
@@
    DEPFusion is proposed, a novel multispectral object detection framework with low-light RGB images enhancement and priority-guided Mamba multi-modality fusion.
@@
    To the best of our knowledge, it is the first time to propose Priority-Guided Serialization with theoretical proof. Based on it, PGMF module is designed to enhance local target modeling and reduce the interference information for multispectral feature fusion.
@@
    The experiment results on DroneVehicle dataset and VEDAI dataset show that DEPFusion performs well on object detection and achieves lightweight computation.
II Related Work
II-1 Multispectral Object Detection
@@
Pixel-wise fusion aggregates two spectral data into a whole images ahead of schedule, and feeds the images to the detector. Cao et.al [15] propose a local–global adaptive dynamic learning method to fuse multi-modal data. To ensure the fusion network is task-optimized, DetFusion [10] employs the detection loss computed from the fused images as guidance during training. Employing input-level fusion of RGB and IR images. Feature-wise fusion focuses on the features from different stages of backbone network. These stage features are fused and fed to the detection head. TSFADet [16] designs an alignment module to predict modality deviations with subsequent feature map calibration. Zhao et.al [17] propose a coarse-to-fine framework to remove redundant information and select features for fusion dynamically. To resolve modality calibration and fusion inaccuracies, $\text{C}^{2}\text{Former}$ [13] introduced an inter-modal cross-attention module, while its Adaptive Feature Sampling module mitigated computational costs from global attention.
@@
II-2 Low-light Images Enhancement
@@
There are many frameworks having been designed for enhancing the low-light images, which lacks enough texture-details due to under exposed characteristic. Brightness enhancement in FourLLIE [18] is achieved by dual-stage estimation of amplitude transformation maps within the frequency domain. DiffLL [19] implements sequential denoising refinements in diffusion models to generate realistic details for low-light image enhancement task, highlighting their practical viability. [20] proposes a framework with guidance of generative perceptual priors obtained from vision-language models, it shows good generalization on real-world data. WaveMamba [21] combines wavelet transformation and Mamba mechanism to achieve low-light images restoration. These works enhance the texture-details of instances in low-light images, which gives important help to the detection task.
Figure 2: The overall architecture of DEPFusion. (a) is the pipeline of DEPFusion, (b) is the architecture of DDE module, while (c) is the processing of CSWM block including the Cross-Scale Scanning Mechanism. (d) shows the Cross-Scale Serialization in the CSWM block.
II-3 Vision Mamba
@@
Mamba [22] mechanism has been paid attention to computer vision field because it can model long-distance dependencies and maintain linear computational complexity. Many vision mamba works has developed and performed well in the vision researches. The visual state-space model VMamba [23] preserves SSM’s global receptive field capability but operates at linear computational cost, enabling order-of-magnitude speed gains. Vim [24] proposes a new vision backbone with bidirectional scanning mamba, This architecture embeds positional cues into sequences while compacting representations through bidirectional state-space compression. Vision mamba is widely adopted for images classification tasks, like SpectralMamba [25], $\text{S}^{2}\text{Mamba}$ [26] etc., and images segmentation tasks like $\text{RS}^{3}\text{Mamba}$ [27], VM-Unet [28] etc. These works could achieve good performance with low computation complexity.
@@
However, most of vision mamba works adopts Z-order serialization for scanning the images, which focuses on global context and suppresses the local dependencies. Therefore, for the remote sensing images with many small objects, the interference information brings degradation to local targets modeling in the multi-modality fusion stage under the Z-order serialization.
III Method
III-A Dual-Domain Mamba Enhancement Module
@@
The low-frequency component in wavelet-domain contains most of information about low-light RGB images, while the high-frequency component only contains small parts of texture-details [21]. Therefore, the low-frequency component could be adopted to enhance the overall brightness of images.
For the texture-details, it is worth noting that frequency domain has good global modeling capacity. The amplitude spectrum contains the energy of details while the phase spectrum contains the structure of the instances in the images. Therefore, these two spectra could be adopted to recover the texture details of images.
@@
Following the above ideas, DDE module is designed for enhancing the low-light RGB images, which is shown in Fig 2. DDE module consists of two parts, the first part is CSWM block, which is designed for low-frequency component enhancement. The second part is FDR block, which is designed to enhance the texture details of images.
III-A1 Cross-Scale Wavelet Mamba Block
@@
where $DWT()$ denotes 2D-DWT operator.
The $LL$ is enhanced by CSWM block, which is designed by adopting the strong context-aware properties of Mamba [22]. To ensure that the local objects dependencies are enhanced, $LL$ is conveyed to three convolution layers with kernel sizes $i\in\left[3,5,7\right]$ to obtain multi-scale feature and global feature:
@@
III-A2 Fourier Details Recovery
The enhanced images $I_{ve}^{n}$ also lacks texture-details after low-frequency enhancement. Therefore, $I_{ve}^{n}$ is conveyed to FDR module for recovering most of texture-details. The structure of FDR module could be seen in Fig 3. FDR contains spectrum recovery networks which could enhance the details of amplitude spectrum (AS) and phase spectrum (PS) obtained by 2D-FFT. The enhanced MS and PS are then transformed into time-domain to obtain rich texture-details images $I_{VE}$. This process could be described as:
@@ III-A2 Fourier Details Recovery
where $SRN()$ denotes the spectrum recovery network, which consists of two DWCONV layers, one ReLU layer, one LayerNorm layer and one Linear layer. $FFT()$ denotes fast 2D-FFT operator, and $IFFT()$ denotes the inverse transformation of 2D-FFT. $I_{VE}$ should be conveyed to CSWM block if it is not in the last level of 2D-DWT.
Algorithm 1 Algorithm of the PGMF Module.
Input: RGB Feature: $\mathcal{F}_{v}$, Infrared Feature: $\mathcal{F}_{i}$
Output: Fusion Feature: $\mathcal{F}_{fus}$
[…]
12:for $t_{v}$, $t_{i}$ in $index_{v}$$index_{i}$ do
13:  $token_{v}\leftarrow\mathcal{F}_{vl}\left[:,:,t_{v}\right]$
14:  $token_{i}\leftarrow\mathcal{F}_{il}\left[:,:,t_{i}\right]$
15:  $Seq_{v}\leftarrow Concat\left(Seq_{v},token_{v}\right)$
16:  $Seq_{i}\leftarrow Concat\left(Seq_{i},token_{i}\right)$
17:end for
[…]
III-B Priority-Guided Mamba Fusion Module
For the feature maps of remote sensing images, there is an obvious difference in the proportion of local target features and background features. However, the Z-order serialization focuses on global modeling, which brings redundant information to weaken the local targets modeling. Therefore, if the scanning prioritizes the feature of local targets, the impact of redundant information is lessened, and the original global feature is retained. The fused feature map with obvious targets details could be obtained by scanning the priority sequences of two modalities. In the later section, we conduct theoretical proof about the advantage of Priority-guided for Mamba modeling.
Following the above ideas, Priority-Guided Mamba Fusion (PGMF) module is designed to achieve high effective RGB-IR feature fusion, which is shown in Fig 4.
III-B1 The Process of PGMF Module
The input feature $\mathcal{F}_{v}$ and $\mathcal{F}_{i}$ are conveyed to small MLPs to reflect into latent space. Moreover, it is worth noting that not only does the feature difference $\varDelta\mathcal{F}$ contain location information of the targets, but also show the regions that be focused mistakenly. Therefore, $\varDelta\mathcal{F}$ is conveyed to Priority Score Network to obtain priority matrices $PMat$. Both of two modalities could obtain their priority features by adding with $PMat$:
@@ III-B1 The Process of PGMF Module
The serialization sets up to obtain the sequences by following the descent priority scores, which separates the target from interference information. Therefore, the scanning starts from the targets and ends up with the background information.
For the reason that the parameter renewing of $A$ matrix of Mamba mechanism relies on the historical information, and it is necessary to improve the contributions of local object feature. The fusion of two sequences is conducted by the format of $Seq$ in Eq (10):
@@
Where $Ser()$ denotes serialization operator and $argSort()$ denotes obtaining the sequences after sorting.
Two sequences are fused together and conveyed to SSM module, which contains forward scanning and backward scanning. The input features of two modalities are added with priority feature by residual connection:
@@
The pseudocode PGMF module could be seen in Algorithm 1.
III-B2 The Theoretical Proof of Priority-Guided Serialization
As known to all, the current output of Mamba [22] depends on the historical information. In addition, the selective learning mechanism is achieved by decaying the hidden state dynamically, which allows Mamba to capture long-range dependencies. Therefore, by proofing that the decaying is related to the position of token, we could proof the advantage of Priority-Guided Serialization, and the fusion way expanded from it. We define the following symbols firstly:
[…]
The hidden state and the output of Mamba could be describe as:
$\begin{split}h_{t}&=\overline{A}h_{t-1}+\overline{B}x_{t}\\ y_{t}&=\overline{C}h_{t}\end{split}$ (12)
where $\overline{A}\in\mathbb{R}^{n\times n},\overline{B}\in\mathbb{R}^{n\times 1},\overline{C}\in\mathbb{R}^{1\times n},x\in\mathbb{R}^{1\times n}$. Therefore, if we define a sequence
$x=\left[x_{1},x_{2},...,x_{t},...,x_{T}\right]$ (13)
the ttht-th output could be described as:
$h_{t}=\sum_{i=0}^{t}{\overline{A}^{i}\overline{B}x_{t-i}}$ (14)
for the contribution of ithi-th token to the ttht-th output, it could be obtained from Eq (14):
$\lVert h_{t}^{i}\rVert=\lVert\left(\prod_{k=i+1}^{t}{A_{k}}\right)B_{i}x_{i}\rVert$ (15)
According to Holder Inequation, the contribution has supremum:
$\lVert h_{t}^{i}\rVert=\lVert\left(\prod_{k=i+1}^{t}{\overline{A}_{k}}\right)\overline{B}_{i}x_{i}\rVert\leq\lVert\prod_{k=i+1}^{t}{\overline{A}_{k}}\rVert\lVert\overline{B}_{i}x_{i}\rVert$ (16)
According to ZOH principle, we could calculate the eigenvalue of $\overline{A}$:
$\lambda\left(\overline{A}_{k}\right)=e^{\varDelta\lambda\left(A\right)_{k}}<1$ (17)
since $\overline{A}=-e^{\ln A}=-A<0$. The spectral radius of $\overline{A}$ satisfies:
$\rho_{k}=\underset{k}{\max}\left[\lambda\left(\overline{A}_{k}\right)\right]<1$ (18)
Following Geland Theorem and Eq (18), we could know that:
$\underset{\left|i-t\right|\rightarrow\infty}{\lim}\lVert\prod_{k=i+1}^{t}{\overline{A}_{k}}\rVert=0$ (19)
The limiting of the item in Eq (16) could be obtained:
$\underset{\left|i-t\right|\rightarrow\infty}{\lim}\lVert\prod_{k=i+1}^{t}{\overline{A}_{k}}\rVert\lVert B_{i}x_{i}\rVert=0$ (20)
For the ithi-th token, if it is far from the ttht-th hidden state, the contribution has infimum:
$\underset{\left|i-t\right|}{\text{inf}}\left(\lVert h_{t}^{i}\rVert\right)=0$ (21)
Following Eq (20), we could obtain the range of the contribution value:
$0\leq\underset{\left(t-i\right)\rightarrow\infty}{\lim}\lVert h_{t}^{i}\rVert\leq\underset{\left(t-i\right)\rightarrow\infty}{\lim}\lVert\prod_{k=i+1}^{t}{A_{k}}\rVert\lVert B_{i}x_{i}\rVert=0$ (22)
Following Squeeze Theorem, we could conclude that:
$\underset{\left(t-i\right)\rightarrow\infty}{\lim}\lVert h_{t}^{i}\rVert=0$ (23)
From Eq (23) we conclude that $A$ matrix decays the information of early tokens and capture new information of late tokens. However, the parameter renewing of $A$ matrix is impacted by the historical information such as the initial tokens. Therefore, for the multispectral fusion sequence, the tokens locating on the head and tail are important. This requirement guides us to construct the fusion sequence by:
$X=x_{vp}+\overline{x_{ip}}$ (24)
where $x_{vp}$ and $x_{ip}$ denote the priority sequences of two modalities. Eq (24) shows that the local target feature are stored in the head and tail of sequence. Under the forward and backward scanning of SSM, the features of two modalities are fused complementarily.
III-B3 The Computational Complexity of PGMF Module
The computational complexity of PSN could be described as:
$O_{psn}=O\left(\sum{\left(H\times W\times K^{2}\times C_{in}\times C_{out}\right)}\right)$ (25)
where $H$,$W$ denote the height and width of images respectively, $K$ denotes the size of convolution kernel, $C_{in}$, $C_{out}$ denote the number input channels and output channels respectively. Notably, $K$, $C_{in}$ and $C_{out}$ are constant parameters. Therefore, Eq (25) could be simplified as:
$O_{psn}=O\left(N\right)$ (26)
where $N$ denotes the number of pixels and $N=H\times W$.
The sorting contributes the computational complexity on the Priority-Guided serialization. Notably, the Radix Sort is adopted by PyTorch, which has linear complexity. For the SSM module, it also has linear complexity:
$\begin{split}O_{pri}&=O\left(N\right)\\ O_{ssm}&=O\left(N\right)\end{split}$ (27)
Therefore, our PGMF module improves the performance of detector and maintains linear complexity, comparing with conventional SSM fusion with Z-order serialization.
III-B4 The Back-Propagation of PGMF Module
In this section, we will derive the gradient of PGMF module during back-propagation. We define the following symbols firstly:
[…]
The output of Mamba [22] $y$ could be described by adopting convolution form:
$y=x\ast\mathcal{K}$ (28)
The sequence fed to SSM module is combined with the priority-guided sub-sequences of two modality. They are obtained from PSN, whose parameters are shared between two branches. Following Eq (28), This process could be described as:
$Y=\sum_{t=0}^{T}{\left\{\mathcal{K}\left(\omega_{t}\right)\left[f_{v}+f_{spn}\left(f_{v}\right)+f_{i}+f_{spn}\left(f_{i}\right)\right]_{t}\right\}}$ (29)
Therefore, the gradient vectors of $f_{v}$$f_{i}$ and $f_{psn}$ could be obtained by the Chain Rule:
$\begin{split}&\nabla\mathcal{L}\left(\theta_{v}\right)=\sum_{t=0}^{T}{\left\{\mathcal{K}\left(w_{t}\right)\left[\nabla f_{v}\left(\theta_{v}\right)+\nabla f_{psn}\left(f_{v}\right)\nabla f_{v}\left(\theta_{v}\right)\right]_{t}\right\}}\\ &\nabla\mathcal{L}\left(\theta_{i}\right)=\sum_{t=0}^{T}{\left\{\mathcal{K}\left(w_{t}\right)\left[\nabla f_{i}\left(\theta_{i}\right)+\nabla f_{psn}\left(f_{i}\right)\nabla f_{i}\left(\theta_{i}\right)\right]_{t}\right\}}\\ &\nabla\mathcal{L}\left(\theta_{psn}\right)=2\sum_{t=0}^{T}{\left\{\mathcal{K}\left(w_{t}\right)\left[\nabla f_{psn}\left(\theta_{psn}\right)\right]_{t}\right\}}\end{split}$ (30)
III-C Loss Function